can you find the points where the line x + 2y = 6 crosses the x axis and y axis?
Answer by nerdybill(7384)

can you find the points where the line x + 2y = 6 crosses the x axis and y axis?
.
To find the point(s) where the function crosses the x-axis (x intercepts) -- set y=0 and solve for x:
x + 2y = 6
x + 2(0) = 6
x + 0 = 6
x = 6
x-intercept is at (6,0)
.
To find the point(s) where the function crosses the y-axis (y intercepts) -- set x=0 and solve for y:
x + 2y = 6
0 + 2y = 6
2y = 6
y = 3
y-intercept is at (0,3)
.
solution: (6,0) and (0,3)